Please Help me with this problem. I am very confused. Determine whether the following equation have a solution or not? √2 x² - 4x – 7 √2 = 0
Answer by nerdybill(7384)

√2 x² - 4x – 7 √2 = 0
.
The coefficients are:
a = √2
b = -4
c = 7√2
.
The determinant (stuff under the radical in the quadratic equation):
b^2 - 4ac
= (-4)^2 - 4(√2)(7√2)
= 16 - 4(2)(7)
= 16 - (8)(7)
= 16 - 56
= -40
.
Since the determinant is negative, you will have two complex solutions.
.
See:
http://www.purplemath.com/modules/quadform3.htm
for additional information